What molal solution is made when adding 90 grams of K2SO4 to 0.55 kg of water?



1
Expert's answer
2021-06-03T05:29:01-0400

Morality is defined as number of moles of solute dissolved in per kg of solvent. 

It is denoted by m.

m = "\frac{number of moles of solute}{weight of solvent in kg}"


Moles of solute = weight of K2SO4"\div" molar mass of K2SO4

Molar mass of K2SO4= 2×39 + 1×32 +4×16 =174g/mol

Moles of K2SO4= "\frac{ 90}{174}= 0.52 mole"


Now 

Molality = "\frac{0.52 mole}{0.55kg }= 0.96 molar"


The solution is 0.96 molar.